OpenStudy (anonymous):

evaluate the following determinant by expanding according to co factors of: (a) the third row (b)the second column 1) |2 -1 2| |1 0 3 | | 0 2 1 |

OpenStudy (anonymous):

i got the answer 7 when i did it but the answers is -7

OpenStudy (anonymous):

(a) 0(-1*3-0*2)-2*(2*3-(2*1)+1(2*0-(-1))=-8+1=-7

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

does the sign change if there is a negative in front or does it remain +-+

OpenStudy (anonymous):

(b) -(-1)*(1*1-0*3) +0*(2*1-0*2)-2(2*3-1*2)=+1-8=-7

OpenStudy (anonymous):

+ -+ -+- +-+ this is how u open, wen there a negative sign in the number the sign automatically gets changed to + or - , if ur multiplying it by - or +. d'u get it ?

OpenStudy (anonymous):

yes thank you for clarifying :)